i m not gettin right ans....


(2.4)^x=(2.6)^x-1


has1. no solution


2. exactly one solution


3. at least two solution


4. infinite number of solution


Grade:12

1 Answers

Askiitians_Expert Yagyadutt
13 years ago

Hii vibhor 

 

(2.4)^x  =  (2.6)^(x-1)

 

(2.4)^x = (2.6)^x / (2.6)

 

(2.4/2.6)^x  = 1/(2.6)

 

(12/13)^x  = 5/13

 

 

So this gives a conclusion that it has one solution lying between  (0,1)

 

 

Regards

Yagya

askiitians_expert